Write the equation for the following graphs: 5. (positive cosine function)
Answer #1

If you need any help in understanding this solution then please ask through comment section. I will be happy to help you.iture yapon) = A cos (Bu +P) +D where mayt min Amplitude A= max-min = 9-43) = 6. De max+ min = 9+(3) - 3 2. phase differende